A scuba fiver dives directly from A to B, and then to C and then to the bottom of the sea at D she then realise that she has only 4 minutes' worth of oxygen left in her tank. She can ascend vertically at 10 cm/s AB=8cm, BC=12m and CD=16m. Can she reach the surface before her oxygen supply runs out?
1
Expert's answer
2019-08-18T12:39:42-0400

The distance in centimeters


"\vert AB \vert= 8 cm, \vert BC \vert= 1200 cm,\vert CD \vert= 1600 cm"


"\vert AD \vert= \vert AB \vert+\vert BC \vert+\vert CD \vert= \\
 =8+1200+1600= 2808 cm"


She can ascend vertically at 10 cm/s.


We will take the time ( T ) to get up:


"T= \frac{\vert AD \vert }{10}=\frac{2808}{10}=280.8 (s)"


"T= 280.8 s"


She has only 4 minutes' worth of oxygen left in her tank.


4 minutes = 4*60 =240 seconds


T> 4 minutes


Can she reach the surface before her oxygen supply runs out?


Answer:

A scuba fiver can't reach the surface before her oxygen supply ends.